# SemNav: A Semantic Segmentation-Driven Approach to Visual Semantic Navigation

### v1

@@ IV-B2 Comparison with the state of the art
We compare here the performance of our SemNav models with state-of-the-art approaches for the ObjectNav problem. We have used the validation set of HM3D [21] dataset. All compared methods use the same experimental evaluation protocol for ObjectNav [16]. The only difference in terms of training data is that our SemNav models use the semantic segmentation information contained in the SemNav dataset.
Model SR (↑) SPL (↑)
DD-PPO [49] 27.9 14.2
OVRL-v2 [50] 64.7 28.1
RRR [51] 30.0 14.0
Frontier based exploration [52, 22] 26.0 15.2
Habitat-Web [5] 57.6 23.8
PirlNav (only IL) [48] (our impl.) 60.9 0.26
PirlNav (only IL) [48] 64.1 0.27
@@
XGX [23] 72.9 0.36
SemNav-OS 72.9 0.34
SemNav-OS+DINO 73 0.34
SemNav-RGBS 74.3 0.35
SemNav-RGBS+DINO 76.2 0.36
SemNav-RGBS+DINO $\rightarrow$ RL 77.75 0.40
@@
TABLE II: Comparison of SemNav approaches with the state-of-the-art models in the ObjectNav task of the HM3D dataset (validation set). We report the performance using the metrics: Succes Rate (SR) and Shortest Path Length (SPL).
@@
Table II includes, to the best of our knowledge, a comparison with the models that define the state of the art in the ObjectNav problem. Based on the results, we would like to highlight the following conclusions. First, it is noteworthy that our SemNav-OS model, the simplest one as it only uses semantic segmentation information from the environment, outperforms all state-of-the-art models except XGX [23]. We believe this underscores the importance of semantic segmentation information for navigation tasks. Second, when incorporating RGB information as a sensor, our SemNav-RGBS model surpasses all other approaches to the problem, setting a new state of the art on the validation set of the HM3D dataset. Third, if we apply finetuning to our best model using RL (SemNav-RGBS+DINO $\rightarrow$ RL), following the strategy in [48], we achieve an additional improvement, bringing the SR to a value above 77%. And fourth, in the qualitative results, we observed an interesting exploratory capability in the SemNav models. Egocentric videos recorded during evaluations in the simulation environment show how SemNav models tend to explore the environment, probing the different rooms they encounter. If a room is not useful for reaching the target object category, they proceed to leave it and continue exploring. The videos show that SemNav models can traverse hallways, inspect rooms, and move between different floors of a building. Moreover, they do not exhibit erratic movements during navigation, reflecting efficient and consistent decision-making. The promising results obtained in the evaluation metrics support these qualitative observations. It is worth mentioning that these behaviors have also been observed in experiments conducted in the real world (see Section IV-C).
@@
Figure 5 illustrates some of these qualitative navigation results. For example, Figure 5 (a) shows how the agent begins its search in a living room, where it does not find the desired object—in this case, a television screen. It then explores all the rooms on the current floor. After failing to locate the object, the agent moves to the upper floor, where it finally identifies the requested object. Figure 5 (b) depicts another interesting navigation episode. Initially, the agent is in a large living room where it cannot detect the requested object category. In the following images, the agent moves through the environment, passing through two bedrooms and viewing the living room from a different perspective. Finally, the agent detects a television and executes a stop action next to it.
@@ IV-C Results in the real world
TABLE III: SR and number of actions per object category, and SDS, reported in real-world experiments.
Figure 6: Comparison of the SR reported in the real world and the simulation environment for the different VSN architectures used.
None of the models were able to reach the bed object, as it was not present in the navigable scene. Interestingly, our SemNav models attempted to climb the stairs when tasked with finding the bed. This behavior arises because, in the simulated environment, the agents can navigate stairs. Moreover, beds are typically located on upper floors, which highlights the type of behavior encoded in the learned navigation policy. Unfortunately, in the real world, our robot lacks this capability. Figure 7 presents some qualitative results.
Figure 7: Qualitative results of the robot successfully navigating in the real world toward a sofa, a television, and a chair.
@@ V Conclusion
Our extensive evaluation shows that SemNav outperforms state-of-the-art ObjectNav models in both the Habitat 2.0 simulator and real-world tests. By leveraging semantic segmentation priors, our model achieves higher success rates and shorter path lengths, even in unseen environments, and exhibits enhanced adaptability in real-world scenarios. The integration of semantic segmentation also reduces the domain gap, a persistent challenge in VSN research.

### v2

@@ IV-B4 Comparison with the state of the art
We compare here the performance of our SemNav models with state-of-the-art approaches for the ObjectNav problem. We have used the official validation set of HM3D [45] dataset. All compared methods use the same experimental evaluation protocol for ObjectNav [42]. The only difference in terms of training data is that our SemNav models use the semantic segmentation information provided in the SemNav dataset.
Model Depth RGB Semantic SR ($\uparrow$) SPL ($\uparrow$)
DD-PPO [40] $\checkmark$ $\checkmark$ $\times$ 27.9 0.14
OVRL-v2 [43] $\times$ $\checkmark$ $\times$ 64.7 0.28
Frontier based exploration [48, 13] $\checkmark$ $\checkmark$ $\times$ 26.0 0.15
Habitat-Web [27] $\checkmark$ $\checkmark$ $\times$ 57.6 0.238
PirlNav (only IL) [26] (our impl.) $\checkmark$ $\checkmark$ $\times$ 60.9 0.26
PirlNav (only IL) [26] $\times$ $\checkmark$ $\times$ 64.1 0.27
@@
XGX [38] $\checkmark$ $\checkmark$ $\times$ 72.9 0.36
RRR [28] $\checkmark$ $\checkmark$ $\times$ 30.0 0.14
Uni-NaVid [51] $\times$ $\checkmark$ $\times$ 73.7 0.37
FiLM-Nav [49] $\checkmark$ $\checkmark$ $\times$ 77.0 0.41
SemNav-OS $\times$ $\times$ $\checkmark$ 72.9 0.34
SemNav-OS+DINO $\times$ $\times$ $\checkmark$ 73 0.34
SemNav-RGBS $\times$ $\checkmark$ $\checkmark$ 74.3 0.35
SemNav-RGBS+DINO $\times$ $\checkmark$ $\checkmark$ 76.2 0.36
SemNav-RGBS+DINO $\rightarrow$ RL $\times$ $\checkmark$ $\checkmark$ 77.75 0.40
@@
TABLE VIII: Comparison of SemNav approaches with the state-of-the-art models in the ObjectNav task of the HM3D dataset (validation set). We report the performance using the metrics: SR and SPL.
@@
Table VIII includes, to the best of our knowledge, a comparison with the models that define the state of the art in the ObjectNav problem. Based on the results, we would like to highlight the following conclusions. First, it is noteworthy that our SemNav-OS model, the simplest one as it only uses semantic segmentation information from the environment, outperforms all state-of-the-art models except XGX [38] and FiLM-Nav [49]. We believe this underscores the importance of semantic segmentation information for navigation tasks. Second, when incorporating RGB information as a sensor, our SemNav-RGBS model surpasses all other approaches in terms of SR, setting a new state of the art on the validation set of the HM3D dataset. Third, if we apply finetuning to our best model using RL (SemNav-RGBS+DINO $\rightarrow$ RL), following the strategy in [26], we achieve an additional improvement, bringing the SR to a value above 77%. Notably, FiLM-Nav [49] achieves a remarkable SPL that exceeds even our RL-finetuned model (SemNav-RGBS+DINO $\rightarrow$ RL); however, FiLM-Nav operates in 2D and is therefore unable to navigate across multiple floors within a building, which limits its applicability in multi-story environments. And fourth, in the qualitative results, we observed an interesting exploratory capability in the SemNav models.
Egocentric videos recorded during evaluations in the simulation environment show how SemNav models tend to explore the environment, probing the different rooms they encounter. If a room is not useful for reaching the target object category, they proceed to leave it and continue exploring. The videos show that SemNav models can traverse hallways, inspect rooms, and move between different floors of a building. Moreover, they do not exhibit erratic movements during navigation, reflecting efficient and consistent decision-making. The promising results obtained in the evaluation metrics support these qualitative observations. It is worth mentioning that these behaviors have also been observed in experiments conducted in the real world (see Section IV-C).
@@
Figure 6 illustrates some of these qualitative navigation results. For example, Figure 6 (a) shows how the agent begins its search in a living room, where it does not find the desired object—in this case, a television screen. It then explores all the rooms on the current floor. After failing to locate the object, the agent moves to the upper floor, where it finally identifies the requested object. Figure 6 (b) depicts another interesting navigation episode. Initially, the agent is in a large living room where it cannot detect the requested object category. In the following images, the agent moves through the environment, passing through two bedrooms and viewing the living room from a different perspective. Finally, the agent detects a television and executes a stop action next to it. We provide more qualitative results, including failure cases, in the following video.
@@ IV-C Results in the real world
The real-world experiments evaluate the performance of the SemNav models in real environments. First, a comparative analysis of different models is conducted, examining the differences in their results. These experiments were carried out in house 1 shown in Figure 4a.
[…]
@@
TABLE IX: SR and number of actions per object category, and SDS, reported in real-world experiments conducted in House 1.
Figure 7: Comparison of the Success Rate (SR) reported in the real world (House 1) and the simulation environment for the different VSN architectures used.
None of the models were able to reach the bed object, as it was not present in the navigable scene of house 1. Interestingly, our SemNav models attempted to climb the stairs when tasked with finding the bed. This behavior arises because, in the simulated environment, the agents can navigate stairs. Moreover, beds are typically located on upper floors, which highlights the type of behavior encoded in the learned navigation policy. Unfortunately, in the real world, our robot lacks this capability. Figure 8 presents some qualitative results in house 1, but more can be found in the following video, including failure cases.
Figure 8: Qualitative results of the robot successfully navigating in the real world House 1 toward a sofa, a television, and a chair.
[…]
Table X shows that, despite navigating in different houses unseen during training, the robot’s performance remains robust, achieving an overall SR of 73.7%, close to the 77.75% obtained in simulation. It is particularly noteworthy that the “bed” category presents failures in two different houses, both caused by collisions with a sofa, as the semantic segmentation used exhibited noise, and both categories are frequently confused by the segmenter. Furthermore, this model exhibits a relatively low number of steps, resulting in more efficient navigation trajectories.
SR / Actions SDS
Chair Bed Toilet Sofa TV Monitor Average
House 1 100% / 29 0% / 61 100% / 99 100% / 40 0% / 39 60% / 53.6 0.018
House 2 100% / 20 100% / 58 100% / 79 100% / 12 100% / 77 100% / 49.2 0.02
House 3 100% / 21 0% / 53 100% / 40 0% / 63 100% / 54 60% / 46.2 0.026
Average 100% / 23.3 33.3% / 57.3 100% / 72.7 66.7% / 38.3 66.7% / 56.7 73.3% / 49.66 0.024
[…]
Figure 10 presents qualitative navigation results in House 2 and House 3. These experiments demonstrate that the agent is capable of successfully navigating toward its target object across different environments, with varying furniture layouts and under different lighting conditions, by leveraging both RGB and semantic segmentation information.
(a) Navigation toward a bed in House 2
(b) Navigation toward a TV monitor in House 3
Figure 10: Qualitative results of the robot successfully navigating in the real world toward a bed in House 2 (left) and toward a TV monitor in House 3 (right), corresponding to the floor plans in Figures 4b and 4c, respectively.
In these real-world experiments, conducted on a TurtleBot 2 platform equipped with an onboard computer featuring an Intel Core i7-10750H CPU @ 2.60 GHz (12 cores) and no dedicated GPU hardware, all models were executed exclusively on the CPU. This setup accurately reflects the computational constraints of typical robotic platforms employed in practical field applications (where no powerful GPU is embedded in the robot). Table XI reports the inference times per step obtained under this configuration. As expected, CPU-based inference leads to a substantial increase in computational latency compared to GPU execution, particularly for the semantic segmentation model. The action decision models (SemNav-OS and SemNav-RGBS) demonstrate relatively lower inference times, yet they remain significantly higher than their GPU-based counterparts. The overall per-step inference time—encompassing both perception and decision stages—constitutes a key limitation for achieving real-time navigation performance in resource-constrained robotic systems. However, these times can be improved by embedding a GPU in the robot, or even using a distributed ROS architecture. For example, the ROS node in charge of semantic segmentation and navigation decisions could have a powerful GPU.
TABLE XI: Temporal breakdown per inference step on CPU (ms). ESANet provides semantic segmentation; SemNav-OS and SemNav-RGBS are action decision models. “Total” = action model + ESANet.
Model Inference Time (ms) + ESANet (ms) Total (ms)
SemNav-OS 79 1400 323
SemNav-RGBS 137 1400 387
@@ V Conclusion
Our extensive evaluation shows that SemNav outperforms state-of-the-art ObjectNav models in both the Habitat 2.0 simulator and real-world tests. By leveraging semantic segmentation priors, our model achieves higher SR and SPL, even in unseen environments, and exhibits enhanced adaptability in real-world scenarios. The integration of semantic segmentation also reduces the domain gap, a persistent challenge in VSN research.
@@
Experimental results show that performance depends on the semantic segmentation step. Models that introduce higher segmentation noise perform worse in simulation than those with lower or no noise, indicating that segmentation quality is a critical factor for SemNav. Obtaining high‑quality semantic segmentations in real environments remains an open challenge, especially in complex or dynamic scenes. Given this challenge, using higher-quality semantic segmentations could further improve SemNav’s performance in real-world scenarios and further reduce the sim-to-real domain gap. In addition, generating these segmentations incurs substantial computational latency, which can limit applicability in real‑time robotic scenarios.
Despite the improved real-world results, real-world navigation necessarily requires a social component that is not addressed in this work. In future work, we plan to integrate social navigation models into the SemNav model to improve interaction with humans and navigation in populated environments. Additionally, future work will explore integrating additional semantic information, such as natural language descriptions or spatial relationships between objects, to enrich the environment representation and improve the agent’s decision-making.